OpenStudy (anonymous):

Write each measure in radians. Express your answer in terms of π. 45 degree

OpenStudy (anonymous):

divide by \(180\) and stick a \(\pi\) next to it

OpenStudy (anonymous):

\[\frac{45\pi}{180}=\frac{\pi}{4}\]

OpenStudy (anonymous):

what about 90 degree? is this same process

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what about 90 degree?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

hello